### v2

# Decomposed Prompting: Probing Multilingual Linguistic Structure Knowledge in Large Language Models

Abstract
Probing the multilingual knowledge of linguistic structure in LLMs, often characterized as sequence labeling, faces challenges with maintaining output templates in current text-to-text prompting strategies. To solve this, we introduce a decomposed prompting approach for sequence labeling tasks. Diverging from the single text-to-text prompt, our prompt method generates for each token of the input sentence an individual prompt which asks for its linguistic label. We test our method on the Universal Dependencies part-of-speech tagging dataset for 38 languages, using both English-centric and multilingual LLMs. Our findings show that decomposed prompting surpasses the iterative prompting baseline in efficacy and efficiency under zero- and few-shot settings. Moreover, our analysis of multilingual performance of English-centric LLMs yields insights into the transferability of linguistic knowledge via multilingual prompting.
Decomposed Prompting: Probing Multilingual Linguistic Structure Knowledge in Large Language Models
Ercong Nie1,2, Shuzhou Yuan3, Bolei Ma2,4, Helmut Schmid1, Michael Färber3, Frauke Kreuter2,4,5 and Hinrich Schütze1,2 1Center for Information and Language Processing (CIS), LMU Munich, 2Munich Center for Machine Learning (MCML), 3ScaDS.AI and TU Dresden, 4Department of Statistics, LMU Munich, 5University of Maryland, College Park nie@cis.lmu.de, shuzhou.yuan@tu-dresden.de, bolei.ma@lmu.de,
1 Introduction
Current Large Language Models (LLMs), such as GPT-3, GPT-4, PaLM, and LLaMA (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023a), have demonstrated remarkable capabilities in in-context learning across a broad spectrum of language understanding and generation tasks (Zhao et al., 2023; Zhang et al., 2023; Ziyu et al., 2023). These models are predominantly trained on massive amounts of English text data, with some limited exposure to other languages. For instance, LLaMA2’s pretraining corpus comprises over 89% English content, with the rest in other languages or code (Touvron et al., 2023b). Yet, these English-centric LLMs 111In this paper, we regard a model pretrained primarily on English text as English-centric. exhibit effective performance in complex multilingual language understanding tasks (Deng et al., 2023; Wang et al., 2023). In multilingual evaluation with prompting, a model performs tasks by directly generating outputs based on a task description and/or a few examples provided in a pivot language (typically English), along with input in a different target language (Ahuja et al., 2023). Despite the remarkable multilingual performance of LLMs, the extent and nature of their cross-lingual capabilities remain underexplored (Ye et al., 2023).
We hypothesize that these models harbor substantial multilingual knowledge. This knowledge, particularly relating to linguistic structure, is commonly conceptualized through sequence tagging tasks (Jurafsky, 2000). However, the current prompting strategies designed for sequence labeling in LLMs are not well suited to test. For instance, behavioral probing methods (Belinkov et al., 2020), aimed at measuring knowledge stored in language models, struggle to adapt to tasks predicting more complex structures. To overcome the challenges in probing the multilingual knowledge of linguistic structure in LLMs characterized as sequence labeling, drawing inspiration from the token-level prompt-based fine-tuning method by Ma et al. (2024), we introduce the decomposed prompting strategy, aiming to probe English-centric LLMs for their understanding of linguistic structure framed as sequence labeling tasks. As shown in Figure 1, instead of employing a single text-to-text prompt for labeling an entire sequence in one step, our method decomposes this process into multiple discrete prompts. More precisely, we first split the input sentence into tokens. Subsequently, we generate an individual prompt for each token which inquires about its linguistic label.
We evaluate our approach on the Universal Dependency (UD) part-of-speech (POS) tagging dataset (Nivre et al., 2020) covering 38 languages with 3 English-centric LLMs and 2 multilingual LLMs. Our approach outperforms the iterative prompting baseline in both zero- and few-shot settings in terms of accuracy and efficiency. Furthermore, our investigation into the multilingual performance of English-centric LLMs offers valuable insights into their capabilities of transferring linguistic knowledge through multilingual prompting.
2 Background and Related Work
Multilinguality of English-Centric LLMs
English-centric LLMs are primarily pretrained on large English text data, with a limited exposion to multilingual data. LLaMA (Touvron et al., 2023a), for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages. LLaMA 2 (Touvron et al., 2023b) expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data. Mistral 7B (Jiang et al., 2023) achieves superior performance and efficiency through the adoption of advanced attention techniques such as Sliding Window Attention (SWA) (Child et al., 2019), facilitating faster inference. To enhance the robustness of multilingual processing, the Byte-level Byte-Pair-Encoding (BBPE) algorithm (Sennrich et al., 2016; Wang et al., 2020) is commonly used for tokenization in LLMs. This approach is able to decompose UTF-8 characters, which are outside the scope of the model vocabulary, into their constituent bytes. Thus, BBPE tokenization equips LLMs with the versatility to handle scripts from any language, theoretically, even those not encountered during training. In summary, limited exposure to non-English data and byte-level encoding capability, these two factors discussed above, jointly contribute to the robust multilingual abilities observed in English-centric LLMs.
Prompting for Sequence Labeling
Prompting LLMs for sequence labeling tasks remains a challenge (Ahuja et al., 2023). While text-to-text prompting is widely adopted across various benchmarking tasks for LLMs (Lai et al., 2023), their application to sequence labeling is hindered by the challenges in maintaining the output templates (Asai et al., 2023). In response, a decent iterative prompting strategy for structured prediction has been introduced (Blevins et al., 2023) (Figure 1). In this approach, the model decodes in step a label for the word at position of the sequence. This predicted label, along with the next word, is then input back into the model to predict the next label. However, the dependency of each token’s prediction on the preceding one substantially slows down the inference process. In contrast, our proposed decomposed prompting method offers improvements in both efficacy and efficiency. Our method is similar to Ma et al. (2024) in that both methods decompose an input sentence into a series of prompts; however, their method is used for fine-tuning, while our method is in an in-context learning paradigm without training.
3 Decomposed Prompting for LLMs
Given a test sequence set , a label set , and an LLM , we approach the task of sequence labeling as follows: for an input sequence of length , , the model is expected to produce a corresponding sequence of labels , where each label is associated with the linguistic feature of the token .
In decomposed prompting, we design a prompt template function which generates a specific prompt for each token. takes the input sequence and an individual token as arguments and returns a prompt for predicting the label of the token. The true label can be optionally included as an argument to ; if included, utilizes to provide a demonstration.
is a sample from the training set. In the few-shot learning scenario, examples in the tuple format are given along with the input sequence , where is a token in , and is the label for . The demonstration of an input sequence is formulated as:
where denotes an optional instruction in natural language, denotes the string concatenation operation. Finally, we use a prompt generator function to create the set of decomposed prompts for an input sequence :
The label of token is predicted as follows:
For each possible label , we obtain the probability that the model predicts this label as the next token and select the most likely label as the predicted label.
4 Experiment and Results
4.1 Experimental Setup
Dataset and Language
We use a subset of the Universal Dependency treebanks (UDPOS) (Nivre et al., 2020) to probe the multilingual linguistic knowledge of LLMs. The UDPOS dataset adopts a universal POS tag set consisting of 17 tags (Appendix A.1.1). Our chosen subset, derived from the XTREME multilingual benchmark (Hu et al., 2020), comprises 38 languages from diverse language families distributions (Appendix A.1.2). We randomly sample 200 instances of each language for the evaluation.
Model and Setup
We experiment on three English-centric LLMs: LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023b), and Mistral-7B (Jiang et al., 2023), as well as two multilingual LLMs: BLOOMZ-7B (Muennighoff et al., 2023) and mTk-Instruct (Wang et al., 2022). All LLMs in our experiment are instruction-tuned versions accessible through the HuggingFace framework (Wolf et al., 2020). We use the weighted average F1 scores for different tags as our evaluation metric. All experiments were conducted on a server with 4 A100-SXM4-80GB GPUs. More details of experimental settings are described in Appendix A.2.
4.2 Overall Results
| Model | Method | Zero-shot | Few-shot | Avg. | ||
| en | mult. | en | mult. | |||
| LLaMA2-7B | Iter | 33.1 | 27.2 | 68.0 | 48.6 | 44.2 |
| Decom | 58.2 | 43.2 | 74.7 | 50.5 | 56.7 | |
| LLaMA2-13B | Iter | 47.6 | 37.4 | 68.0 | 52.6 | 51.4 |
| Decom | 67.3 | 54.7 | 77.3 | 54.5 | 63.5 | |
| Mistral-7B | Iter | 65.2 | 54.3 | 80.2 | 58.9 | 64.7 |
| Decom | 63.6 | 61.8 | 85.0 | 64.4 | 68.7 | |
| BLOOMZ-7B | Decom | 20.6 | 17.6 | 44.1 | 36.2 | 29.6 |
| mTK-Instruct | Decom | 47.6 | 43.1 | 57.3 | 44.7 | 48.2 |
We evaluate the performance of iterative prompting, the baseline method, and decomposed prompting, our proposed method, for English and multilingual POS tag labeling tasks under zero- and few-shot settings. The few-shot examples and the prompts employed in our experiment are presented in Appendix B.2 for reference. Our preliminary experiment to explore the influence of the number of few-shot samples () reveals a mild impact on performance once increases to around 10. More details are provided in Appendix C.
Superiority in Efficacy
The overall results for English-centric LLMs, as detailed in Table 1, demonstrate that our proposed decomposed prompting obviously outperforms the iterative prompting baseline across both zero- and few-shot settings, in both English and multilingual evaluations. This trend holds true for all three English-centric models tested, with the sole exception in the zero-shot setting for the English evaluation with the Mistral-7B model, where Decom slightly lags behind Iter (63.6 vs. 65.2). In addition, English-centric LLMs outperform multilingual LLMs by a considerable margin. The complete experimental results are displayed in Appendix D.
| BLOOMZ-7B | LLaMA2-7B | Mistral-7B | Avg. | |
| zero-shot | ||||
| few-shot |
Superiority in Efficiency
In addition to superior performance, decomposed prompting offers enhanced efficiency during inference, especially in few-shot prompting. As demonstrated in Table 2, our proposed method achieves, on average, a 2.4-fold increase in speed compared to the baseline in the zero-shot prompting setting and a 6.7-fold increase in the few-shot setting. The efficiency advantage is less obvious with Mistral, owing to Mistral’s implementation of a modified attention mechanism designed to enhance inference efficiency.
5 Multilingual Analysis
Figure 2 provides a stratified view of decomposed prompting performance by language family and script, under both zero- and few-shot settings on the Mistral model. The results indicate that Indo-European languages generally achieve higher F1 scores compared to their non-Indo-European counterparts. Notably, the presence of few-shot examples consistently improves the overall performance across all categories, but the box plot also shows that some languages are negatively impacted by the use of English demonstrations. As discussed in §2, English-centric LLMs are adept at tokenizing words from Latin or Cyrillic scripts into subtokens. For scripts less familiar to these models, they often default to breaking down the text into UTF-8 encodings, which may lead to suboptimal representations for languages using these less common scripts. Thus, to capture a more nuanced understanding of LLM performance across linguistic varieties, we categorize languages not only by family but also by script type. Figure 2(b) illustrates that, in both few-shot and zero-shot settings, languages with known scripts tend to yield better performance than unknown scripts. An exception to this trend is observed among the language group with smaller corpora in the zero-shot setting.
To further understand the impact of English demonstrations on languages with varied properties in multilingual prompting, we delve deeper into the cross-lingual transferability of English-centric LLMs and conduct a detailed analysis of individual language performance. We begin by quantifying the linguistic proximity of each tested language to English. This was achieved by calculating the cosine similarity between language vectors (Littell et al., 2017) that incorporate syntactic, phylogenetic, and geographic attributes, among others, following Nie et al. (2023) and Ma et al. (2023). Further information on the computation of language similarity is available in Appendix A.3. From Figure 3, we observe that the performance gain from few-shot prompting is more substantial for languages that are linguistically closer to English, as indicated by the upward trend on the right side of the plot. Remarkably, languages distant from English may even experience a decline in performance when using English demonstrations.
6 Conclusion
In conclusion, we propose decomposed prompting, a simple yet effective prompting method specially designed for sequence labeling tasks, addressing the difficulties of LLM benchmarking on sequence labeling tasks. Our method outperforms iterative prompting techniques in terms of accuracy and efficiency in different experimental settings. By applying decomposed prompting to UDPOS dataset, we probe the multilingual linguistic structure knowledge of English-centric LLMs. Our multilingual investigation reveals that gain from few-shot decomposed prompting is generally more pronounced for languages closer to English.
Limitations
Although our proposed decomposed prompting method achieves overall remarkable performance in terms of both accuracy and efficiency, it has limitations for some special cases, for example, it can not well handle the case where the same word occurs twice in a sentence with different POS tags. Besides, the efficiency of decomposed prompting suffers as the length of the input sequence and the complexity of the task increase. Our study uses decomposed prompting methods for part-of-speech (POS) tagging as a means to evaluate the multilingual structural knowledge of English-centric Large Language Models (LLMs). This provides a foundational assessment of the models’ capabilities. Nevertheless, extending the application scope of this methodology to probe more intricate aspects of linguistic structure is necessary. Future research could beneficially apply decomposed prompting to the analysis of complex linguistic phenomena, including sentence chunking and named entity recognition, to gain a deeper understanding of the nuanced capabilities of LLMs in processing and understanding language.
Acknowledgements
We want to thank the anonymous reviewers for their valuable feedback. This work was supported by the German Research Foundation (DFG) under grant SCHU 2246/14-1, Munich Center for Machine Learning (MCML), and China Scholarship Council (CSC).
Appendix A Experimental Setup Details
Details of the experimental setup are introduced in this section.
A.1 Dataset and Languages
A.1.1 POS Tag Set
Figure 4 shows the pos tag set in UD. We also use the text in the box as the task instruction in our experiments.
A.1.2 Profile of Languages
As Figure 5 shows, our experiment involves 38 languages with diverse language family distributions.
A.2 Baselines and Settings
Iterative Prompting (Iter)
Blevins et al. (2023) introduced a structured prompting approach that iteratively labels an entire sentence by appending each predicted label to the context along with the subsequent word. This method is employed as a strong baseline in our study.
Decomposed Prompting (Decom)
To evaluate our proposed approach, we employ the prompt template outlined in §3 to decompose the entire sequence into a set of individual prompts for prediction. In our experiments, we use the 17 POS tags themselves as the label words, i.e., we expect the model to directly predict a tag from the tagset shown in Figure 4 by selecting the tag with the highest logit.
Zero- and Few-Shot Prompting
We devised two experimental scenarios for multilingual prompting—zero-shot and few-shot—to evaluate the performance of both approaches under different conditions. In the zero-shot setting, only an English task instruction is provided alongside the input in the target language. The text in Figure 4, which outlines the tag set information, serves as the instruction in our experiments. In few-shot prompting, we supplement the prompt with a few English demonstrations, structured according to the prompt template of each method. For Decom, we randomly select an example for each tag type from the English training set to create a demonstration. For a fair comparison, the same number of demonstrations are used for the Iter baseline.
A.3 Language Similarity Computation
Malaviya et al. (2017) and Littell et al. (2017) proposed lang2vec, language vectors to represent various linguistic features for languages. A language can be represented by five vectors, containing syntactic, phonological, phonetic, phylogenetic, and geographical features, respectively. Linguistic similarities among different languages with respect to these linguistic features can be calculated through the cosine similarity. In our study, we utilized the language vectors provided by lang2vec to calculate the cosine similarity between target languages and English. We used a rank-based similarity score to average the rank of languages in each feature dimension. Table 3 illustrates the computation details.
| syn. | syn_rank | pho. | pho_rank | inv. | inv_rank | fam. | fam_rank | geo. | geo_rank | rank_score | |
| eng-nld | 92.43 | 37 | 81.83 | 18 | 76.28 | 36 | 44.51 | 35 | 99.96 | 37 | 32.6 |
| eng-deu | 90.26 | 36 | 80.60 | 15 | 78.68 | 37 | 54.49 | 37 | 99.76 | 35 | 32.0 |
| eng-ukr | 84.73 | 32 | 85.83 | 32 | 74.91 | 33 | 15.03 | 30 | 99.28 | 26 | 30.6 |
| eng-por | 84.24 | 31 | 90.46 | 35 | 74.03 | 28 | 10.14 | 22 | 99.68 | 33 | 29.8 |
| eng-ell | 78.31 | 25 | 95.35 | 37 | 74.74 | 32 | 15.03 | 32 | 98.96 | 22 | 29.6 |
| eng-pol | 78.64 | 26 | 85.83 | 29 | 74.09 | 29 | 15.03 | 31 | 99.63 | 32 | 29.4 |
| eng-bul | 85.78 | 35 | 85.83 | 30 | 74.38 | 30 | 13.73 | 27 | 99.01 | 23 | 29.0 |
| eng-ita | 85.78 | 34 | 85.83 | 28 | 72.94 | 26 | 11.21 | 23 | 99.53 | 30 | 28.2 |
| eng-rus | 81.18 | 29 | 85.83 | 31 | 74.63 | 31 | 16.80 | 33 | 95.81 | 17 | 28.2 |
| eng-ron | 79.60 | 27 | 90.46 | 34 | 73.42 | 27 | 11.89 | 24 | 99.22 | 25 | 27.4 |
| eng-spa | 82.16 | 30 | 85.83 | 27 | 72.83 | 25 | 9.71 | 21 | 99.59 | 31 | 26.8 |
| eng-lit | 69.33 | 18 | 80.42 | 14 | 75.58 | 34 | 19.39 | 34 | 99.44 | 27 | 25.4 |
| eng-afr | 84.94 | 33 | 81.83 | 17 | 75.91 | 35 | 50.46 | 36 | 86.84 | 6 | 25.4 |
| eng-fra | 81.18 | 28 | 75.28 | 7 | 72.24 | 24 | 9.71 | 20 | 99.93 | 36 | 23.0 |
| eng-est | 77.35 | 24 | 85.83 | 25 | 70.81 | 19 | 0.23 | 15 | 99.45 | 28 | 22.2 |
| eng-hun | 69.40 | 19 | 85.83 | 24 | 70.66 | 18 | 0.33 | 18 | 99.46 | 29 | 21.6 |
| eng-fin | 71.08 | 21 | 87.05 | 33 | 70.00 | 17 | 0.19 | 13 | 99.19 | 24 | 21.6 |
| eng-eus | 62.36 | 13 | 85.29 | 21 | 70.00 | 16 | 3.33 | 19 | 99.76 | 34 | 20.6 |
| eng-urd | 61.63 | 12 | 85.83 | 26 | 71.98 | 23 | 12.71 | 25 | 92.54 | 13 | 19.8 |
| eng-mar | 56.50 | 8 | 80.42 | 13 | 71.57 | 22 | 13.73 | 28 | 89.80 | 11 | 16.4 |
| eng-wol | 63.92 | 14 | 85.83 | 23 | 69.73 | 15 | 0.17 | 10 | 96.24 | 18 | 16.0 |
| eng-hin | 61.63 | 11 | 78.35 | 10 | 70.91 | 20 | 12.71 | 26 | 91.10 | 12 | 15.8 |
| eng-fas | 50.03 | 3 | 78.35 | 11 | 70.94 | 21 | 13.73 | 29 | 94.23 | 14 | 15.6 |
| eng-ind | 72.66 | 22 | 90.92 | 36 | 67.09 | 12 | 0.12 | 4 | 79.16 | 1 | 15.0 |
| eng-heb | 75.15 | 23 | 72.55 | 5 | 69.10 | 14 | 0.13 | 6 | 97.16 | 20 | 13.6 |
| eng-ara | 65.11 | 16 | 70.09 | 3 | 68.38 | 13 | 0.15 | 9 | 97.04 | 19 | 12.0 |
| eng-tur | 50.68 | 4 | 81.83 | 16 | 67.09 | 11 | 0.14 | 7 | 98.25 | 21 | 11.8 |
| eng-zho | 71.08 | 20 | 72.55 | 4 | 66.94 | 10 | 0.33 | 16 | 88.42 | 9 | 11.8 |
| eng-kaz | 44.77 | 1 | 83.64 | 19 | 66.59 | 9 | 0.14 | 8 | 95.22 | 16 | 10.6 |
| eng-vie | 66.04 | 17 | 78.35 | 9 | 65.81 | 8 | 0.19 | 11 | 85.25 | 3 | 9.6 |
| eng-tel | 52.07 | 6 | 80.42 | 12 | 64.76 | 4 | 0.19 | 14 | 89.18 | 10 | 9.2 |
| eng-tgl | 60.89 | 10 | 85.83 | 22 | 64.76 | 5 | 0.13 | 5 | 82.15 | 2 | 8.8 |
| eng-tam | 51.36 | 5 | 85.29 | 20 | 64.37 | 3 | 0.11 | 3 | 87.95 | 8 | 7.8 |
| eng-kor | 55.29 | 7 | 74.65 | 6 | 63.83 | 2 | 0.33 | 17 | 86.93 | 7 | 7.8 |
| eng-tha | 63.95 | 15 | 78.35 | 8 | 65.40 | 7 | 0.11 | 2 | 85.25 | 4 | 7.2 |
| eng-yor | 60.04 | 9 | 66.77 | 2 | 65.29 | 6 | 0.10 | 1 | 94.98 | 15 | 6.6 |
| eng-jpn | 50.03 | 2 | 66.77 | 1 | 56.88 | 1 | 0.19 | 12 | 85.65 | 5 | 4.2 |
Appendix B More Details of Decomposed Prompting Method
B.1 Intuition
This method draws inspiration from the step-by-step thinking process humans employ when annotating linguistic features within a sentence. Typically, humans approach such tasks incrementally, addressing each token individually. Mirroring this intuitive strategy, our method first decomposes an input sentence into tokens. Subsequently, we generate a distinct prompt for each token, thereby transforming the sequence labeling task into a series of focused, manageable prompts. Figure 6 illustrates the generation of sequence labeling prompts for the German sentence “Viel Erfolg!” via decomposed prompting.
Figure 6 illustrates the generation of sequence labeling prompts for the German sentence “Viel Erfolg!” via decomposed prompting.
An example of a template function is illustrated as follows.
B.2 Prompt Details
Zero- and few-shot prompts used in this work are shown in Figure 8 (decomposed prompting) and Figure 9 (iterative prompting).
Appendix C Few-Shot Ablation Study
we investigate the impact of the number of few-shot examples on the performance in the decomposed prompting. We randomly select 8 languages (en, de, el, fa, hi, hl, ru, zh) and explore their performance dynamics with the increasing of the few-shot samples. Figure 7 shows that overall, when is small, increasing the number of samples bring performance improvement. As continues to increase, the performance tends to be stable and even gets worse when samples are too many.
Appendix D Full Results
Full experimental results are displayed in Table 4 (Mistral 7B), Table 5 (LLaMA2 7B), Table 6 (LLaMA 13B), Table 7 (BLOOMZ 7B), Table 8 (mTk 13B), and Table 9 (few-shot ablation study).
| language | en | af | ar | bg | de | el | es | et | eu | fa | fi | fr | he | |
| zero-shot | Iter | 65.2 | 67.8 | 57.2 | 68.6 | 65.0 | 55.0 | 64.8 | 49.4 | 35.6 | 58.3 | 50.2 | 65.4 | 51.5 |
| Decom (prob.) | 63.6 | 66.0 | 67.8 | 74.4 | 68.6 | 62.7 | 68.6 | 58.0 | 54.1 | 68.5 | 60.2 | 63.5 | 66.4 | |
| Decom (gen.) | 45.3 | 43.8 | 49.6 | 50.5 | 49.0 | 50.7 | 43.3 | 53.6 | 50.7 | 56.0 | 55.5 | 40.5 | 55.6 | |
| few-shot | Iter | 80.2 | 66.4 | 65.0 | 77.3 | 66.9 | 56.4 | 70.8 | 53.7 | 50.7 | 57.4 | 63.9 | 67.7 | 66.4 |
| Decom (prob.) | 85.0 | 76.9 | 48.1 | 82.4 | 78.3 | 52.3 | 82.7 | 65.2 | 48.8 | 57.3 | 64.4 | 76.9 | 66.6 | |
| Decom (gen.) | 81.4 | 74.8 | 44.3 | 80.4 | 77.0 | 46.3 | 82.0 | 64.0 | 48.1 | 54.1 | 63.6 | 76.4 | 64.9 | |
| Decom (prob.) + I | 83.4 | 77.9 | 42.4 | 76.9 | 77.8 | 33.6 | 77.6 | 64.6 | 57.4 | 42.9 | 67.6 | 74.8 | 58.5 | |
| Decom (gen.) + I | 78.7 | 75.8 | 34.0 | 74.9 | 76.6 | 24.7 | 76.4 | 62.6 | 56.8 | 34.4 | 64.5 | 73.4 | 54.5 | |
| language | hi | hu | id | it | ja | kk | ko | lt | mr | nl | pl | pt | ro | |
| zero-shot | Iter | 61.3 | 50.6 | 54.7 | 64.0 | 42.2 | 36.7 | 39.9 | 52.8 | 39.1 | 60.4 | 66.5 | 63.9 | 66.2 |
| Decom (prob.) | 37.1 | 58.6 | 61.0 | 68.6 | 56.3 | 57.8 | 47.4 | 68.2 | 61.0 | 69.4 | 73.5 | 68.4 | 68.5 | |
| Decom (gen.) | 35.6 | 46.7 | 41.8 | 45.1 | 48.9 | 50.2 | 42.2 | 60.3 | 56.7 | 46.8 | 59.5 | 43.1 | 44.6 | |
| few-shot | Iter | 65.7 | 50.4 | 70.0 | 67.2 | 42.0 | 43.8 | 42.6 | 63.2 | 54.4 | 66.6 | 70.9 | 75.1 | 65.9 |
| Decom (prob.) | 67.8 | 71.3 | 73.9 | 76.2 | 59.8 | 50.0 | 44.0 | 67.5 | 48.9 | 80.6 | 78.6 | 77.8 | 77.8 | |
| Decom (gen.) | 66.2 | 70.8 | 73.0 | 76.0 | 57.1 | 50.2 | 43.4 | 67.1 | 48.9 | 77.2 | 78.3 | 76.9 | 77.0 | |
| Decom (prob.) + I | 57.6 | 66.5 | 70.4 | 72.2 | 54.2 | 58.4 | 49.2 | 69.9 | 53.1 | 78.5 | 76.7 | 75.0 | 76.4 | |
| Decom (gen.) + I | 55.3 | 63.9 | 68.2 | 70.3 | 53.1 | 57.9 | 48.2 | 69.5 | 52.7 | 76.9 | 75.7 | 74.2 | 75.1 | |
| language | ru | ta | te | th | tl | tr | uk | ur | vi | wo | yo | zh | avg. | |
| zero-shot | Iter | 68.2 | 39.2 | 51.1 | 54.1 | 65.0 | 47.7 | 67.0 | 56.0 | 41.7 | 31.5 | 41.3 | 58.8 | 54.3 |
| Decom (prob.) | 74.4 | 55.2 | 63.8 | 63.0 | 62.9 | 55.2 | 74.1 | 54.2 | 59.9 | 39.6 | 49.7 | 59.2 | 61.8 | |
| Decom (gen.) | 54.7 | 52.2 | 57.4 | 50.1 | 51.3 | 43.2 | 57.4 | 40.3 | 45.9 | 29.2 | 43.3 | 55.7 | 48.7 | |
| few-shot | Iter | 74.0 | 52.0 | 62.4 | 57.1 | 37.3 | 62.0 | 68.2 | 59.6 | 41.0 | 25.2 | 39.0 | 62.3 | 58.9 |
| Decom (prob.) | 79.9 | 37.5 | 61.4 | 58.2 | 73.4 | 62.7 | 77.7 | 51.3 | 52.6 | 42.0 | 47.8 | 65.8 | 64.4 | |
| Decom (gen.) | 78.0 | 33.9 | 61.3 | 56.9 | 73.4 | 62.6 | 76.2 | 45.7 | 52.8 | 42.0 | 47.6 | 64.5 | 63.0 | |
| Decom (prob.) + I | 76.8 | 35.7 | 67.0 | 45.8 | 74.9 | 63.7 | 75.1 | 40.5 | 59.4 | 43.1 | 49.2 | 62.9 | 62.3 | |
| Decom (gen.) + I | 73.9 | 28.0 | 66.6 | 42.9 | 74.9 | 62.6 | 73.4 | 32.9 | 59.7 | 43.2 | 48.6 | 61.4 | 59.9 | |
| language | en | af | ar | bg | de | el | es | et | eu | fa | fi | fr | he | |
| zero-shot | Iter | 33.1 | 38.8 | 30.2 | 33.2 | 34.5 | 38.1 | 38.9 | 19.7 | 11.8 | 17.7 | 26.0 | 37.5 | 21.3 |
| Decom (prob.) | 58.2 | 45.1 | 49.6 | 55.9 | 53.3 | 50.4 | 44.7 | 37.7 | 36.4 | 40.5 | 41.3 | 46.8 | 39.5 | |
| Decom (gen.) | 53.8 | 46.8 | 38.5 | 45.8 | 57.1 | 54.3 | 52.4 | 28.6 | 20.2 | 35.9 | 39.8 | 53.1 | 37.5 | |
| few-shot | Iter | 68.0 | 56.1 | 58.0 | 63.4 | 56.9 | 48.7 | 55.3 | 46.5 | 41.3 | 51.1 | 50.5 | 54.2 | 54.0 |
| Decom (prob.) | 74.7 | 60.0 | 29.9 | 64.7 | 63.0 | 30.6 | 55.7 | 53.0 | 44.4 | 29.7 | 62.9 | 54.4 | 42.8 | |
| Decom (gen.) | 62.1 | 51.0 | 25.7 | 60.3 | 52.4 | 23.9 | 50.3 | 48.3 | 42.9 | 26.0 | 56.8 | 49.5 | 37.5 | |
| Decom (prob.) + I | 68.2 | 55.9 | 23.7 | 61.6 | 61.0 | 20.2 | 52.5 | 43.2 | 40.8 | 22.7 | 49.4 | 54.8 | 35.4 | |
| Decom (gen.) + I | 63.4 | 53.2 | 19.0 | 57.9 | 56.2 | 12.0 | 47.8 | 39.3 | 40.0 | 15.5 | 46.4 | 51.2 | 30.1 | |
| language | hi | hu | id | it | ja | kk | ko | lt | mr | nl | pl | pt | ro | |
| zero-shot | Iter | 35.2 | 29.3 | 31.1 | 35.1 | 28.7 | 13.6 | 19.8 | 24.9 | 13.2 | 37.5 | 37.7 | 38.4 | 32.0 |
| Decom (prob.) | 36.9 | 47.0 | 46.9 | 46.7 | 32.4 | 39.0 | 29.0 | 34.9 | 45.3 | 54.9 | 54.0 | 48.6 | 43.6 | |
| Decom (gen.) | 34.8 | 47.4 | 39.1 | 45.2 | 30.9 | 33.0 | 33.2 | 37.7 | 42.0 | 51.1 | 44.1 | 48.5 | 42.6 | |
| few-shot | Iter | 54.0 | 41.0 | 51.3 | 49.6 | 40.0 | 43.2 | 25.0 | 52.5 | 50.3 | 52.2 | 52.4 | 52.0 | 53.8 |
| Decom (prob.) | 45.8 | 62.6 | 60.9 | 56.4 | 40.2 | 51.4 | 48.2 | 56.3 | 47.3 | 58.9 | 67.2 | 60.3 | 63.6 | |
| Decom (gen.) | 42.4 | 57.0 | 56.5 | 51.6 | 34.1 | 47.5 | 44.7 | 51.7 | 43.5 | 51.3 | 64.2 | 54.5 | 55.5 | |
| Decom (prob.) + I | 30.6 | 52.3 | 54.1 | 51.3 | 37.3 | 46.6 | 41.9 | 46.5 | 45.7 | 64.2 | 65.4 | 55.2 | 56.4 | |
| Decom (gen.) + I | 24.1 | 50.6 | 49.5 | 44.1 | 32.9 | 46.0 | 40.7 | 45.3 | 34.5 | 60.2 | 62.0 | 51.2 | 51.8 | |
| language | ru | ta | te | th | tl | tr | uk | ur | vi | wo | yo | zh | avg. | |
| zero-shot | Iter | 29.8 | 19.2 | 13.8 | 29.2 | 28.6 | 22.2 | 30.3 | 20.7 | 29.7 | 13.3 | 13.7 | 32.2 | 27.2 |
| Decom (prob.) | 55.8 | 38.0 | 34.0 | 37.5 | 57.3 | 48.3 | 57.4 | 31.6 | 39.5 | 27.6 | 29.1 | 42.9 | 43.2 | |
| Decom (gen.) | 48.7 | 25.5 | 36.9 | 34.6 | 66.3 | 45.9 | 48.8 | 28.4 | 35.3 | 18.7 | 21.8 | 44.0 | 40.4 | |
| few-shot | Iter | 58.2 | 30.9 | 54.3 | 49.4 | 37.3 | 34.4 | 57.7 | 44.0 | 46.5 | 40.7 | 39.3 | 52.0 | 48.6 |
| Decom (prob.) | 67.2 | 31.7 | 44.7 | 36.5 | 46.8 | 58.1 | 62.9 | 27.1 | 41.4 | 39.9 | 37.1 | 64.8 | 50.5 | |
| Decom (gen.) | 62.3 | 25.3 | 43.5 | 34.7 | 45.4 | 55.9 | 59.4 | 23.7 | 40.7 | 36.2 | 35.5 | 50.9 | 45.8 | |
| Decom (prob.) + I | 59.6 | 20.3 | 38.4 | 20.9 | 63.1 | 54.1 | 59.9 | 19.3 | 49.7 | 32.2 | 33.8 | 48.2 | 45.1 | |
| Decom (gen.) + I | 56.9 | 12.5 | 34.5 | 16.7 | 58.8 | 52.7 | 57.5 | 13.0 | 47.8 | 29.7 | 31.7 | 44.2 | 41.0 | |
| language | en | af | ar | bg | de | el | es | et | eu | fa | fi | fr | he | |
| zero-shot | Iter | 47.6 | 37.4 | 43.2 | 44.5 | 45.7 | 38.4 | 46.8 | 37.0 | 26.5 | 42.0 | 40.7 | 45.5 | 40.0 |
| Decom (prob.) | 67.3 | 60.1 | 54.4 | 62.7 | 63.6 | 60.5 | 55.9 | 49.9 | 37.4 | 59.8 | 62.6 | 53.4 | 55.4 | |
| Decom (gen.) | 59.2 | 54.1 | 45.0 | 52.5 | 57.5 | 51.3 | 56.3 | 37.6 | 36.7 | 49.7 | 50.2 | 54.7 | 44.3 | |
| few-shot | Iter | 68.0 | 62.3 | 57.4 | 69.9 | 60.3 | 57.9 | 66.7 | 44.8 | 41.0 | 49.1 | 54.2 | 63.2 | 59.8 |
| Decom (prob.) | 77.3 | 67.8 | 33.2 | 67.6 | 67.5 | 35.0 | 62.6 | 58.5 | 46.9 | 34.7 | 62.8 | 64.8 | 48.4 | |
| Decom (gen.) | 65.3 | 59.1 | 25.1 | 61.3 | 58.6 | 24.6 | 53.5 | 51.8 | 45.8 | 27.4 | 55.4 | 55.9 | 43.9 | |
| Decom (prob.) + I | 74.3 | 67.6 | 25.9 | 60.7 | 70.5 | 21.5 | 59.1 | 51.4 | 44.1 | 21.8 | 59.1 | 63.1 | 40.3 | |
| Decom (gen.) + I | 68.7 | 64.4 | 19.2 | 58.7 | 66.2 | 12.4 | 53.9 | 47.9 | 42.2 | 15.5 | 54.0 | 59.7 | 35.0 | |
| language | hi | hu | id | it | ja | kk | ko | lt | mr | nl | pl | pt | ro | |
| zero-shot | Iter | 45.0 | 38.8 | 40.9 | 41.8 | 42.8 | 24.1 | 29.8 | 41.2 | 30.5 | 36.6 | 42.2 | 43.3 | 43.1 |
| Decom (prob.) | 53.8 | 57.6 | 57.4 | 54.8 | 48.3 | 51.8 | 45.1 | 54.3 | 50.2 | 62.0 | 66.4 | 56.6 | 57.9 | |
| Decom (gen.) | 45.4 | 47.9 | 48.2 | 51.3 | 35.9 | 48.7 | 35.3 | 43.2 | 48.7 | 56.9 | 58.2 | 51.3 | 51.4 | |
| few-shot | Iter | 51.6 | 46.1 | 60.8 | 62.7 | 46.5 | 32.0 | 26.6 | 50.8 | 52.7 | 61.0 | 64.4 | 68.9 | 58.9 |
| Decom (prob.) | 45.4 | 69.8 | 62.2 | 61.2 | 44.6 | 52.3 | 46.1 | 63.0 | 49.6 | 65.4 | 68.1 | 62.3 | 63.6 | |
| Decom (gen.) | 37.3 | 60.5 | 55.8 | 54.5 | 40.7 | 49.4 | 42.6 | 58.4 | 46.9 | 54.9 | 61.4 | 54.3 | 54.9 | |
| Decom (prob.) + I | 31.4 | 64.2 | 55.3 | 55.3 | 38.1 | 51.7 | 47.1 | 58.9 | 52.5 | 65.4 | 60.2 | 56.3 | 60.4 | |
| Decom (gen.) + I | 23.4 | 60.0 | 50.2 | 52.4 | 35.5 | 49.0 | 45.3 | 56.9 | 50.8 | 61.1 | 58.2 | 54.1 | 56.1 | |
| language | ru | ta | te | th | tl | tr | uk | ur | vi | wo | yo | zh | avg. | |
| zero-shot | Iter | 42.6 | 21.8 | 22.5 | 45.6 | 29.3 | 29.9 | 39.8 | 35.1 | 36.0 | 24.4 | 24.1 | 45.2 | 37.4 |
| Decom (prob.) | 66.5 | 49.1 | 50.8 | 44.6 | 66.5 | 56.9 | 65.7 | 47.2 | 45.3 | 34.5 | 47.7 | 58.7 | 54.7 | |
| Decom (gen.) | 55.2 | 46.2 | 54.1 | 44.2 | 73.1 | 52.8 | 57.3 | 40.2 | 45.4 | 29.9 | 39.6 | 52.5 | 48.7 | |
| few-shot | Iter | 64.9 | 33.5 | 51.5 | 51.5 | 60.2 | 46.3 | 61.6 | 45.4 | 41.8 | 36.3 | 31.6 | 52.1 | 52.6 |
| Decom (prob.) | 71.0 | 30.4 | 54.4 | 40.1 | 74.0 | 54.1 | 69.0 | 30.1 | 47.5 | 39.4 | 36.2 | 66.6 | 54.5 | |
| Decom (gen.) | 63.3 | 21.9 | 51.3 | 33.9 | 70.9 | 52.2 | 61.4 | 22.1 | 45.2 | 38.1 | 34.8 | 56.5 | 48.3 | |
| Decom (prob.) + I | 63.3 | 22.3 | 52.2 | 23.5 | 70.7 | 53.9 | 62.4 | 19.0 | 48.4 | 36.9 | 36.4 | 56.7 | 49.4 | |
| Decom (gen.) + I | 59.8 | 14.1 | 48.4 | 18.5 | 70.2 | 53.2 | 59.1 | 12.0 | 47.1 | 34.5 | 34.5 | 52.7 | 45.6 | |
| language | en | af | ar | bg | de | el | es | et | eu | fa | fi | fr | he | |
| zero-shot | Iter | 6.4 | 7.2 | 10.9 | 7.6 | 9.5 | 8.4 | 8.2 | 12.4 | 7.5 | 7.3 | 9.3 | 9.0 | 9.6 |
| Decom (prob.) | 20.6 | 20.5 | 14.5 | 19.7 | 26.2 | 18.3 | 18.2 | 22.3 | 19.0 | 12.8 | 19.2 | 19.4 | 15.2 | |
| Decom (gen.) | 28.7 | 18.3 | 16.4 | 22.6 | 26.8 | 22.7 | 24.9 | 21.2 | 25.0 | 11.3 | 20.9 | 20.9 | 21.8 | |
| few-shot | Iter | 30.9 | 6.4 | 14.4 | 23.8 | 19.3 | 7.7 | 23.2 | 16.6 | 28.4 | 11.1 | 22.3 | 25.1 | 7.5 |
| Decom (prob.) | 44.1 | 33.1 | 28.7 | 35.9 | 44.0 | 39.2 | 33.6 | 39.0 | 38.4 | 25.6 | 38.5 | 35.6 | 34.3 | |
| Decom (gen.) | 40.6 | 31.0 | 25.5 | 31.4 | 39.5 | 35.8 | 30.5 | 36.9 | 33.8 | 21.6 | 36.8 | 31.0 | 33.6 | |
| Decom (prob.) + I | 33.3 | 24.7 | 27.2 | 35.2 | 30.0 | 31.0 | 30.1 | 36.5 | 37.4 | 24.7 | 34.4 | 29.0 | 29.2 | |
| Decom (gen.) + I | 33.3 | 24.5 | 27.1 | 35.0 | 29.7 | 30.4 | 30.0 | 36.4 | 37.1 | 24.5 | 34.5 | 28.9 | 29.1 | |
| language | hi | hu | id | it | ja | kk | ko | lt | mr | nl | pl | pt | ro | |
| zero-shot | Iter | 3.9 | 13.0 | 10.0 | 9.1 | 2.8 | 4.5 | 8.5 | 7.8 | 0.4 | 9.1 | 9.9 | 8.6 | 8.8 |
| Decom (prob.) | 12.0 | 27.0 | 17.7 | 23.1 | 13.5 | 17.7 | 19.5 | 23.6 | 12.4 | 18.6 | 23.6 | 19.5 | 19.6 | |
| Decom (gen.) | 15.2 | 21.9 | 17.3 | 26.2 | 26.2 | 16.8 | 21.3 | 23.4 | 25.8 | 14.7 | 23.2 | 27.8 | 24.3 | |
| few-shot | Iter | 20.5 | 13.4 | 30.5 | 19.0 | 6.3 | 17.0 | 5.9 | 15.0 | 35.2 | 20.8 | 17.9 | 27.4 | 13.4 |
| Decom (prob.) | 27.0 | 38.2 | 43.8 | 33.9 | 25.9 | 45.6 | 35.0 | 40.3 | 39.6 | 39.8 | 39.7 | 34.4 | 33.3 | |
| Decom (gen.) | 24.8 | 36.9 | 41.2 | 31.1 | 22.5 | 43.8 | 32.7 | 39.5 | 28.0 | 36.5 | 36.5 | 31.7 | 32.0 | |
| Decom (prob.) + I | 25.6 | 32.3 | 36.0 | 30.7 | 25.3 | 45.2 | 27.7 | 41.0 | 44.5 | 29.0 | 34.7 | 30.4 | 32.5 | |
| Decom (gen.) + I | 25.6 | 32.2 | 35.9 | 30.6 | 25.1 | 45.1 | 27.7 | 41.0 | 43.7 | 28.6 | 34.6 | 30.3 | 32.5 | |
| language | ru | ta | te | th | tl | tr | uk | ur | vi | wo | yo | zh | avg. | |
| zero-shot | Iter | 6.8 | 5.0 | 5.1 | 6.8 | 3.9 | 9.0 | 5.2 | 6.6 | 4.2 | 1.4 | 7.2 | 7.6 | 7.4 |
| Decom (prob.) | 26.1 | 15.0 | 7.9 | 8.7 | 7.8 | 15.5 | 23.7 | 8.1 | 14.4 | 11.0 | 18.9 | 21.7 | 17.6 | |
| Decom (gen.) | 27.9 | 20.7 | 12.8 | 2.7 | 1.9 | 17.4 | 28.1 | 12.8 | 25.7 | 21.1 | 28.3 | 26.0 | 20.6 | |
| few-shot | Iter | 20.3 | 24.3 | 47.0 | 3.1 | 22.5 | 20.9 | 20.9 | 15.5 | 18.3 | 16.5 | 16.9 | 20.7 | 18.8 |
| Decom (prob.) | 41.9 | 36.5 | 48.2 | 25.0 | 41.9 | 37.9 | 39.6 | 26.2 | 26.9 | 34.1 | 39.2 | 40.8 | 36.2 | |
| Decom (gen.) | 36.8 | 33.5 | 41.7 | 23.1 | 41.9 | 36.4 | 37.0 | 24.7 | 24.5 | 33.2 | 36.5 | 35.7 | 33.2 | |
| Decom (prob.) + I | 37.0 | 34.1 | 39.0 | 13.7 | 57.8 | 38.0 | 35.8 | 26.4 | 34.0 | 30.3 | 33.3 | 32.8 | 32.9 | |
| Decom (gen.) + I | 36.9 | 33.9 | 38.8 | 13.6 | 57.8 | 38.0 | 35.4 | 26.4 | 33.9 | 30.3 | 33.3 | 32.6 | 32.7 | |
| language | en | af | ar | bg | de | el | es | et | eu | fa | fi | fr | he | |
| zero-shot | Decom (gen.) | 47.6 | 45.7 | 37.8 | 48.9 | 48.9 | 45.8 | 40.0 | 45.3 | 41.5 | 44.2 | 46.8 | 42.6 | 42.6 |
| few-shot | Decom (gen.) | 49.0 | 41.0 | 16.2 | 37.6 | 43.9 | 31.0 | 37.2 | 34.8 | 33.9 | 33.4 | 32.1 | 38.5 | 34.1 |
| Decom (gen.) + I | 57.3 | 51.9 | 27.4 | 47.2 | 55.4 | 40.1 | 50.1 | 41.2 | 43.6 | 48.1 | 42.4 | 49.9 | 45.6 | |
| language | hi | hu | id | it | ja | kk | ko | lt | mr | nl | pl | pt | ro | |
| zero-shot | Decom (gen.) | 40.6 | 38.7 | 39.3 | 39.3 | 32.9 | 46.1 | 29.2 | 47.4 | 47.5 | 42.8 | 46.1 | 40.6 | 49.4 |
| few-shot | Decom (gen.) | 23.8 | 33.5 | 39.9 | 36.5 | 14.3 | 32.4 | 17.7 | 37.5 | 34.9 | 42.7 | 36.1 | 37.1 | 35.6 |
| Decom (gen.) + I | 44.7 | 36.2 | 51.9 | 45.7 | 44.6 | 45.7 | 26.7 | 45.7 | 48.8 | 55.3 | 46.2 | 48.9 | 51.5 | |
| language | ru | ta | te | th | tl | tr | uk | ur | vi | wo | yo | zh | avg. | |
| zero-shot | Decom (gen.) | 45.9 | 39.4 | 51.3 | 47.1 | 59.3 | 46.9 | 47.4 | 37.9 | 48.4 | 22.3 | 37.5 | 42.8 | 43.1 |
| few-shot | Decom (gen.) | 33.5 | 28.1 | 50.9 | 21.9 | 65.7 | 34.7 | 31.2 | 17.7 | 33.9 | 10.5 | 22.4 | 17.2 | 32.5 |
| Decom (gen.) + I | 43.8 | 38.0 | 55.3 | 46.6 | 70.5 | 46.0 | 41.5 | 36.0 | 49.0 | 19.8 | 38.6 | 34.5 | 44.7 | |
| en | de | el | fa | hi | nl | ru | zh | avg. | |
| 0 | 63.6 | 68.6 | 62.7 | 68.5 | 37.1 | 69.4 | 74.4 | 59.2 | 62.9 |
| 1 | 72.4 | 76.1 | 69.2 | 65.9 | 62.5 | 75.1 | 69.4 | 60.2 | 68.9 |
| 3 | 77.0 | 75.8 | 70.9 | 70.7 | 69.7 | 71.5 | 75.2 | 70.1 | 72.6 |
| 5 | 84.4 | 80.9 | 76.7 | 75.3 | 70.9 | 80.4 | 81.8 | 65.9 | 77.0 |
| 7 | 86.2 | 79.8 | 76.5 | 79.3 | 71.4 | 78.8 | 83.2 | 68.3 | 77.9 |
| 9 | 85.4 | 78.5 | 72.8 | 78.2 | 73.1 | 77.6 | 81.8 | 66.4 | 76.7 |
| 11 | 84.9 | 80.2 | 73.9 | 78.6 | 72.9 | 81.3 | 82.3 | 65.7 | 77.5 |
| 13 | 84.1 | 78.7 | 73.1 | 77.9 | 72.6 | 79.1 | 80.7 | 65.4 | 76.5 |
| 15 | 86.2 | 80.0 | 73.6 | 71.6 | 73.7 | 82.7 | 83.4 | 67.9 | 77.4 |
| 17 | 85.8 | 78.3 | 72.9 | 70.1 | 73.0 | 80.9 | 80.8 | 66.6 | 76.1 |